OpenStudy (anonymous):

Triangle PST is a 45-45-90 degree triangle with vertices S(4, -3), T(-2,3) and m

OpenStudy (rizags):

how did you get that distance?

OpenStudy (rizags):

i got it nevermind lol ok im working on it

OpenStudy (rizags):

my distance formula system of equations is not yielding a pretty answer lemme try something else

OpenStudy (rizags):

did you use the distance formula to find TP?

OpenStudy (rizags):

I GOT IT!

OpenStudy (rizags):

and its NOOOOT a raidcal

OpenStudy (rizags):

are you still hereEEE?

OpenStudy (rizags):

HELOOOOOOO>!???!?? i SOLVed It

OpenStudy (anonymous):

Oh sorry! I got disconnected! D:

OpenStudy (rizags):

ok are you here now?

OpenStudy (rizags):

tell me if you know what the distance formula is

OpenStudy (anonymous):

Okay the distance formula is \[d=\sqrt{(x _{1}-x _{2})+(y _{1}-y _{2)}}\]

OpenStudy (rizags):

so, we know the points S(4, -3), T(-2,3), and we know that \[SP=6\sqrt{2}\] and \[TP=12\]

OpenStudy (rizags):

therefore, we can let the point at P be (x,y)

OpenStudy (rizags):

using the distance formula, one can create a system of two equations

OpenStudy (rizags):

that look like so: \[6\sqrt{2}=\sqrt{(x-4)^{2}+(y-(-3))^{2}} and 12=\sqrt{(x-(-2))^{2}+(y-3))^{2}} \]

OpenStudy (anonymous):

And then we just solve the system for x and y?

OpenStudy (rizags):

the way I solved it was by looking at the second equation and thinking that all that stuff inside the radical has to be equal to 144, in order for the entire radical to be equal to 12. Furthermore, the only way this can occur is if x=10, and y=3. As it turns out this also works for the first equation!! the answer must be (10,3). Here is the wolfram alpha proof of solution, but actually full out solving these systems by hand would be incredibly tedious http://www.wolframalpha.com/input/?i=solve+12%3Dsqrt%28%28x%2B2%29%5E2%2B%28y-3%29%5E2%29%2C+6sqrt%282%29%3Dsqrt%28%28x-4%29%5E2%2B%28y%2B3%29%5E2%29%2C+for+x+and+y

OpenStudy (rizags):

by the way, that first solution given by wolfram alpha is invalidated by the question's given that THE point P is in QUADRANT 1, which means that the x and y values of the coordinate MUST both be positive!!!

OpenStudy (rizags):

so yea, i think thats the answeR!!

OpenStudy (anonymous):

Thank you so so so much for this! I was having such a hard time with it. I checked the back of the book, and (10,3) is correct!

OpenStudy (rizags):

niiiiice!

OpenStudy (rizags):

kk fanned u lol